S2 —TPC24 (in the air section of S2, rebuild in 2017)
FRS Home Page -
–
S2 –
Pictures –
Drawings –
Cable Connections – TPC Electronics – Memo for preparation before beam time
Present Configuration (as of 2020/October/22)
- Installed in the air section of S2.
- Gap (vertical) = 8 cm (This means -3200 V for the drift field voltage)
- Normal orientation.
- Now grid is behind TPC.
- Calibration positions should be x = -12, 0, 12 mm and y = -6, 0, +6 mm.
- Maximum allowed HV: -3200 V (drift), +1100 V (anode), -700 V (calibration, PMT).
For heavy ions, anode and PMT voltages should be much lower than the maximum values.
Status during previous beam time
2021/Feb/15-Feb/22 (12C beam, S533/BARB + EXPERT test)
- HV(anode) = +1050V, HV(drift) = -3200V, HV(grid-pmt) = -700V
- Analog raw signal around 150--200 mV.
- Calibration file is /d/land4/sec/s533/tpc24_calib_0026.lmd (Logbook p41-43)
- Parameter sets used during experiment is found
here.
2020/Apr/08--Apr/27 (Pb beam, S468, S482, S481)
- Applied voltages: anode HV = +600 V, drift = -2400 V, grid-PMT = -350 V.
2020/Feb/28--Mar/16 (Xe beam, S474, S459, S480)
- Applied voltages: anode HV = +600 V, drift = -3200 V, grid-PMT = -400 V
- Issue: 3rd anode had very week signal in TDC. This could be due to amplifier or detector itself.
- Calibration
pictures and
parameters (with mask trigger) in run0110 (Feb.29)
- To have consistent track from all 4 TPCs, one could add the values on page 5 of
this memo to the grid position
- Y position is now caculated properly even with different trigger (time reference is subtracted).
2020/Feb/20--Feb/26 (Kr beam, Ca fragments, Cave-C S467)
- Applied voltages: anode HV = +700 V, drift HV = -3200V, Fiber-PMT = -500 V
- Issue: 3rd anode had very week signal in TDC. This could be due to amplifier or detector itself.
- Calibration
pictures and
parameters (with mask trigger) in run0090
- Y position with other triggers (particularly SC81) should be calculated by taking into account time references.
--
NisanYaprakBenal - 2026-10-30